OpenStudy (anonymous):

f(x)=x^2 g(x)=x-3 3) g(f(-2))

OpenStudy (bahrom7893):

this time it's g(f(-2), so first find g(f(x)

OpenStudy (bahrom7893):

g(f(x)) = (x^2) - 3 g(f(-2)) = 4 - 3 = 1

OpenStudy (anonymous):

-7